OpenStudy (anonymous):

define -t(x) =x^3 - 3x^2 + 2x -3 problem -t(-1)

jimthompson5910 (jim_thompson5910):

wherever you see an x, replace it with a -1 t(x) =x^3 - 3x^2 + 2x -3 t(-1) =(-1)^3 - 3(-1)^2 + 2(-1) -3 Then evaluate t(-1) =(-1)^3 - 3(-1)^2 + 2(-1) -3 t(-1) =-1 - 3(1) + 2(-1) -3 t(-1) =-1 - 3 - 2 - 3 t(-1) = - 9

OpenStudy (anonymous):

oh thank you , thank you!!!

jimthompson5910 (jim_thompson5910):

np